OpenStudy (anonymous):

Identify a factor of this trinomial: 36x^2-5xy-24y^2

myininaya (myininaya):

\[36(-24)=12(3)(-4)(6)=-48(18)=16(-54)=32(-27)\]

myininaya (myininaya):

\[36(24)=-32(27) ; -5xy=-32xy+27xy\] \[36x^2-32xy+27xy-24y^2\]

myininaya (myininaya):

factor by grouping

myininaya (myininaya):

\[4x(9x-8y)+3y(9x-8y)\]

myininaya (myininaya):

i leave the last step to you

OpenStudy (anonymous):

alright so it'll be (4x+3y)(9x-8y) Thanks

myininaya (myininaya):

yep